Table 1 Stock Prices and Multiples for the Footwear Industry,January 2006 Ticker Name Stock Price ($) Market Cap ($ Enterprise Value P/E Price/BookEnterprise Enterprise millions) (millions) Value/Sales 1.43 2.19 |value/EBITDA 84.2 21,830 5,088 58.723,514 1,257 800 683 497 373 230 106 20,518 16.643.59 ,593 14.995.02 3,451 14.912.41 2.71 1.91 2.02 1.87 367 13.322.29 22611.97 1.75 NKE PMMAY Puma AG RBKReebok WWWWolverine World Wide BWS 8.75 9.02 8.58 9.53 9.09 6.88 9.28 7.44 6.66 7.55 10.75 Nike 312.05 22.1 43.36 17.09 13.7 30.05 19.9 19.96 6.83 10.4 1,25317.42 1,01922.62 61417.63 524 20.72 0.47 0.62 0.89 1.48 1.06 2328.661.120.92 0.87 0.76 Brown Shoe Sketchers Stride Rite SRR DECKDeckers Outdoor WEYSWeco Group RCKY İTooky Shoes & Boots DFZR.G. Barry Corp BOOT LaCross Footwear ! 8.11 1.28 75 12.09
Suppose that in January 2006, Kenneth Cole Productions had sales of S518 million, EBITDA of S55.6 million, excess cash of $100 million, $3 million of debt, and 21 million shares outstanding 518.00 55.60 100.00 3.00 Sales EBITDA Cash Debt Shares outstandin,g 21 a. Using the average enterprise value to EBITDA multiple in Table 1, estimate KCP's Price per share b. What range of share prices do you estimate based on the highest and lowest enterprise value to EBITDA multiples in Table 1? Highest price estimate Lowest price estimate c. Using the average enterprise value to sales multiple in Table 1, estimate KCP's share Price per share d. What range of share prices do you estimate based on the highest and lowest enterprise value to sales multiples in Table 1? Highest price estimate Lowest orice estimate
Answer #1

Answer a:

We calculate below average enterprise value to EBIDTA multiple:

Enterprise Value (S millions Stock Price Market Cap (S Enterprise Price/BookValue/ Sales valuelt Enterprise Name P/E millions

Given:

KCP's EBITDA = $55.6 million

Debt = $3 million

Excess Cash = $100 million

Shares outstanding = 21 million

KCP's Enterprise value (EV) = KCP's EBITDA * Average enterprise value to EBIDTA multiple = 55.6 * 8.48583 = $471.81 million

KCP's Market capitalization = EV - Debt + Excess cash = 471.81 - 3 + 100 = $568.81 million

Price per share = Market capitalization /Shares outstanding = 568.81 / 21 = $27.09

Price per share $27.09

Answer b:

Highest Average enterprise value to EBIDTA multiple = 10.75

Lowest Average enterprise value to EBIDTA multiple = 6.6

Highest share value = (55.6 *10.75 - 3 +100) / 21 = $33.08

Lowest share value = (55.6 * 6.66 - 3 +100) / 21 = $22.25

Highest price estimate $33.08 Lowest price estimate $22.25

Answer c:

As calculated above (in answer a):

Average enterprise value to sales multiple = 1.06583

KCP's sales =518 million

Price per share = (518 *1.06583 - 3 + 100) / 21 = $30.91

Price per share $30.91

Answer d:

Highest enterprise value to sales multiple = 2.19

Lowest enterprise value to sales multiple = 0.47

Highest share value = (518 * 2.19 - 3 +100) / 21 = $58.64

Lowest share value = (518 * 0.47- 3 +100) / 21 = $16.21

Highest price estimate $58.64 Lowest price estimate $16.21